# Conceptualization; Methodology; Resources; Investigation; Writing original draft; Writing, review, and editing

### v1

@@ 3.4 GAN-based methods
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (2)
@@
Similarly, for CI metal artifacts reduction also, a conditional generative adversarial networks (cGAN) were proposed by Wang et al. [63]. The approach involves training a cGAN to learn mapping from artifact-affected CTs to artifact-free CTs. During inference, the cGAN generated CT images with removed artifacts. Additionally, a band-wise normalization method was proposed as a preprocessing step to improve the performance of cGAN. The method was evaluated on post-implantation CTs recipients, and the quality of the artifact-corrected images was quantitatively assessed using P2PE. The results demonstrate promising artifact reduction, outperforming the previously proposed techniques. The authors evaluates the quality of artifact-corrected images using a quantitative metric based on segmentations of intracochlear anatomical structures. Specifically, the segmentation results obtained from a previously published method were compared between real preimplantation CTs and artifact-corrected CTs generated by the proposed method. The ASE was used as a metric to assess the accuracy of the segmentation. The paper reports that the proposed method achieves an ASE of 0.18 mm, which is approximately half of the error obtained with a previously proposed technique.
3.5 RNN/LSTM-based methods
@@
$\begin{split}A_{f}&=\sigma(W_{f}*L_{i}+U_{f}*L_{j-1}+b_{f}),\quad A_{i}=\sigma(W_{i}*L_{i}+U_{i}*L_{j-1}+b_{i}),\\ L_{j}&=A_{j}\odot\tanh(V_{c}),\\ A_{j}&=\sigma(W_{0}*L_{i}+U_{0}*L_{j-1}+b_{j}),\\ V_{c}&=A_{f}*V_{c-1}+A_{i}*\text{cell\_state}(W_{c}*L_{i}+U_{c}*L_{j-1}+b_{c}).\end{split}$ (4)
@@
In Figure 9, "oz," "cz," "t5," and "t6" refer to specific electrode placements or positions on the scalp in the international 10-20 system for electroencephalography (EEG) or event-related potential (ERP) recordings. These positions represent specific areas on the scalp where electrodes are attached to measure electrical activity in the brain. The amplitude represents the intensity or strength of the electrical signal detected at a particular point on the scalp, reflecting the neural activity in the corresponding brain region. The parameters "N75," "P100," and "N145" refer to specific components or peaks of ERPs obtained from EEG recordings. ERPs are electrical responses recorded from the brain in response to specific stimuli or events, and they reflect the neural processing associated with those stimuli. Besides, $I/P$ represent inputs, $O/P$ for output, and $W$ represent Weights.
3.6 AE-based methods
@@
As a point of the case, the scientific paper [16] delves into the pivotal objective of enhancing speech perception for CI users in noisy conditions, recognizing the critical role of noise reduction (NR) in this pursuit. The proposed method, named DDAE-NR, has been proven effective in restoring clean speech. The study focuses on evaluating the DDAE-based NR using envelope-based vocoded speech, mimicking CI devices. The procedure of DDAE-based NR can be split into two main stages as shown in Figure 10: training and testing. During the training phase, a collection of pairs of noisy and clean speech signals is prepared. These signals are initially transformed into the frequency domain using an fast Fourier transform (FFT). The logarithmic amplitudes of the noisy and clean speech spectra are then used as inputs and outputs, respectively, for the DDAE model.
Key findings underscore the superior intelligibility of DDAE-based NR in vocoded speech compared to state-of-the-art (SOTA) conventional methods, indicating its potential implementation in CI speech processors. However, the study acknowledges the use of noise-vocoded speech simulation for evaluation and emphasizes the need for further validation with real CI recipients in clinical settings, addressing potential inconsistencies in the transition to actual CI devices.
Figure 10: The configuration of a DDAE-based NR system for enhancing speech perception [16].
@@
[…]
The proposed min-max similarity (MMS) methodology in [61] represents a groundbreaking approach to semi-supervised segmentation networks, particularly in the context of medical applications such as endoscopy surgical tool segmentation and CI surgery. MMS is introduced through dual-view training with contrastive learning, utilizing classifiers and projectors to create negative, positive, and negative pairs. The inclusion of pixel-wise contrastive loss ensures the consistency of unlabeled predictions. In the evaluation phase, MMS was tested on four public endoscopy surgical tool segmentation datasets and a manually annotated CI surgery dataset. The results demonstrate its superiority over state-of-the-art semi-supervised and fully supervised segmentation algorithms, both quantitatively and qualitatively. Notably, MMS exhibited successful recognition of unknown surgical tools, providing reliable predictions, and achieved real-time video segmentation with an impressive inference speed of about 40 frames per second. This signifies the potential of MMS as a highly effective and efficient tool in medical image segmentation, showcasing its applicability in real-world surgical scenarios.
@@ 3.7 RL-based methods
4 Applications of DL-based medical CI
This section explores the application of deep learning in the field of cochlear implants, encompassing tasks such as speech denoising and enhancement, segmentation for precise identification and analysis of cochlear structures, thresholding, imaging, localization of CI, and more. Figure 13 provides a comprehensive overview of AI-based applications for CIs and theirs associated benefits. Furthermore, Table 5 summarizes various applications based on AI techniques, highlighting their performance, and pros and cons.
@@ 4 Applications of DL-based medical CI
4.1 Speech denoising and enhancement
@@
Table 5: Summary of the performance and limitations of specific DL applications dedicated to CIs. In cases where multiple tests are conducted, only the best performance is reported.
Ref.
Year
DLM
Description
BP (%)
Limitations
PLA
[82]
2021
NC and DDAE_T
The DL model comprises Siren noise at 6dB, a classifier, and the DDAE. The transfer learning (TL) is incorporated to help reduce the number of parameters in the model.
PESQ= 3.037
@@
The performance of the DL model can be degraded when there is a mismatch between the training and testing data.
No
[22]
2023
DeepACE
Takes the raw audio signal as input and generates a denoised electrodogram, which represents the electrical stimulation patterns applied to the electrodes over time.
STOI =0.807
Heavy data reliance impedes performance with insufficient or biased data, posing generalization challenges. Computational complexity, interpretability issues, and vulnerability are concerns.
Yes [^15] https://github.com/APGDHZ/DeepACE2.0
[29]
2021
U-Net
The proposed approach involves using deep virtual sensing to estimate microphone signals, improving speech quality in cocktail party scenarios for hearing aid and CI users.
PESQ=2, STOI=0.53, SI-SDR=-28.91dB
Challenges include long delay impracticality for hearing aids, binaural cues omission, insufficient network optimization, and exploring robustness in high-reverberation environments.
No
[83]
2022
CNN
A web-based research platform called Nautilus, which utilizes DL techniques for automated image processing in cochlear implantation-related studies.
DCS= 80 ±$\pm 3$%, ASD= 0.19 ±$\pm 0.04$mm
Dependence on large amounts of annotated training data poses a challenge. Ensuring generalizability across various imaging conditions and populations is crucial.
Yes161616nautilus_info@oticonmedical.com available upon reasonable request
[60]
2023
AE+Res-CA
Is a vestibule segmentation network for CT images. It is based on the basic encoder-decoder framework and incorporates Res-CA Block, GCPFE and ACE-Loss Function
DCS=94.77 ±$\pm 2.45$%, ASD=0.06 mm
Limited training data, small object handling, generalization capability, and computational resource constraints are areas of concern.
No
[84]
2019
cGAN
Is a DL-based method for accurately localizing CI electrode contacts in CT images, facilitating customization of CI settings.
success rate of localization =96.7%
coarse resolution, metal artifacts, difficulty isolating contacts from cortical bones, site-dependent image quality, and variability among electrode array manufacturers.
No
[32]
2022
DRL
Is a novel method for accurately extracting ROIs in the inner ear from CT scans. The approach achieves high precision and demonstrates promising results for surgical planning in cochlear implantation.
IoU=0.84
@@
Small dataset, image quality variability, and computational requirements not specified.
No
[75]
2020
U-Net
Is a multi-scale DL framework for automatic segmentation and measurement of the human cochlea in clinical ultra-high-resolution CT images, with potential applications in personalized CI surgery.
DSC=0.90 ±$\pm 0.03$
@@
The work include a small dataset, potential variability in image quality, lack of external validation, and limited assessment of clinical utility and computational requirements.
No
[85]
2023
UNETR
Explores the feasibility of using a DL method based on the UNETR model for automatic segmentation of the cochlea in temporal bone CT images.
DSC=0.92
The small dataset used, the variability in image quality, and the absence of specifications regarding computational requirements.
No
[86]
2019
3D U-NET
Is a two-level training approach using a DL method to accurately segment the intra-cochlear anatomy in head CT scans. The method combines an active shape model-based method and a 3D U-Net model
DSC=0.87
Limited dataset, variable image quality, lack of external validation, limited assessment of clinical utility, and no specifications on computational requirements.
No
[34]
2023
AlexNet
The study assesses repeatability, thresholds, and tonotopic patterns using a DL-based algorithm, providing insights into inner ear function and potential clinical applications.
Acc=83.8%
Potential dependence on the quality of the input data, limited generalizability to different patient populations or implant systems, and the need for further external validation and comparison with expert visual inspection.
No
@@ 4.2 Imaging
4.3 Segmentation
DL, ML, and AI have revolutionized CI segmentation, enabling precise identification and analysis of cochlear structures in various imaging modalities. Li et al. [85] applied a UNETR model to automatically segment cochlear structures in temporal bone CT images, enhancing surgical planning and cochlear implantation outcomes. Reda et al. [93] developed an automatic segmentation method for intra-cochlear anatomy in post-implantation CT scans, facilitating the customization of sound processing strategies for individual CI recipients. Moudgalya et al. [94] employed a modified V-Net CNN to segment cochlear compartments in $\mu$CT images, enabling precise quantification of local drug delivery for potential treatment of sensorineural hearing loss. Wang et al. [95] focused on metal artifact reduction and intra-cochlear anatomy segmentation in CT images using a multi-resolution multi-task deep network, benefiting CI recipients. Heutink et al. [75] developed a DL framework for the automatic segmentation and analysis of cochlear structures in ultra-high-resolution CT images, providing accurate measurements for personalized implant planning in cochlear surgery. Zhang et al. [86] utilized a 3D U-Net DL method to achieve accurate segmentation of intra-cochlear anatomy in head CT images, facilitating optimal programming of CIs and improving hearing outcomes. These studies highlight the significant impact of DL, ML, and AI in advancing CI segmentation, ultimately leading to improved patient care and treatment outcomes.
4.4 Thresholding
@@
4.5 Localization of CI
DL methods have been instrumental in CI localization applications, providing accurate and automated solutions. Chi et al. [84] proposed a DL-based method for precise localization of electrode contacts in CT images. Their approach utilized cGANs to generate likelihood maps, which were then processed to estimate the exact location of each contact. Radutoiu et al. [32] focused on the automatic extraction of ROIs in full head CT scans of the inner ear. By leveraging AI, they achieved high precision in ROI localization, facilitating accurate surgical planning for insertion. Noble et al. [98] and Zhao et al. [99, 100], developed AI-based systems to automatically identify and position electrode arrays in CT images. These technologies enable large-scale analyses of the relationship between electrode placement and hearing outcomes, leading to potential advancements in implant design and surgical techniques. Heutink et al. [75] employed DL for the automatic segmentation and localization of the cochlea in ultra-high-resolution CT images. This approach allows for precise measurements that can be used in personalized planning, reducing the risk of intracochlear trauma and optimizing surgical outcomes. These studies showcase the significant contributions of DL and AI in localization applications, enabling accurate and efficient identification, positioning, and analysis of electrode arrays and facilitating improved surgical planning and outcomes.
4.6 Other

### v3

@@ III-D GAN-based methods
$\begin{split}\min_{G}\max_{D}V(D,G)=&\,\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]\\ &+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]\end{split}$ (2)
@@
Similarly, for CI metal artifacts reduction also, a conditional generative adversarial networks (cGAN) were proposed by Wang et al. [71]. The approach involves training a cGAN to learn mapping from artifact-affected CTs to artifact-free CTs. During inference, the cGAN generated CT images with removed artifacts. Additionally, a band-wise normalization method was proposed as a preprocessing step to improve the performance of cGAN. The method was evaluated on post-implantation CTs recipients, and the quality of the artifact-corrected images was quantitatively assessed using P2PE. The results demonstrate promising artifact reduction, outperforming the previously proposed techniques. The authors evaluates the quality of artifact-corrected images using a quantitative metric based on segmentations of intracochlear anatomical structures. Specifically, the segmentation results obtained from a previously published method were compared between real preimplantation CTs and artifact-corrected CTs generated by the proposed method. The ASE was used as a metric to assess the accuracy of the segmentation. The paper reports that the proposed method achieves an ASE of 0.18 mm, which is approximately half of the error obtained with a previously proposed technique. Gogate et al. [79] propose a robust real-time audio-visual speech enhancement framework for CIs. By leveraging a GAN and DNN, the framework effectively addresses visual and acoustic speech noise in real-world environments. Experimental results demonstrate significant improvements in speech quality and intelligibility, offering potential benefits for CI users in noisy social settings.
III-E RNN/LSTM-based methods
@@
$\begin{split}A_{f}&=\sigma(W_{f}*L_{i}+U_{f}*L_{j-1}+b_{f}),\\ A_{i}&=\sigma(W_{i}*L_{i}+U_{i}*L_{j-1}+b_{i}),\\ L_{j}&=A_{j}\odot\tanh(V_{c}),\\ A_{j}&=\sigma(W_{0}*L_{i}+U_{0}*L_{j-1}+b_{j}),\\ V_{c}&=A_{f}*V_{c-1}+A_{i}*\text{cell\_state}(W_{c}*L_{i}+U_{c}*L_{j-1}+b_{c}).\end{split}$ (4)
@@
In Figure 9, ”oz,” ”cz,” ”t5,” and ”t6” refer to specific electrode placements or positions on the scalp in the international 10-20 system for electroencephalography (EEG) or event-related potential (ERP) recordings. These positions represent specific areas on the scalp where electrodes are attached to measure electrical activity in the brain. The amplitude represents the intensity or strength of the electrical signal detected at a particular point on the scalp, reflecting the neural activity in the corresponding brain region. The parameters ”N75,” ”P100,” and ”N145” refer to specific components or peaks of ERPs obtained from EEG recordings. ERPs are electrical responses recorded from the brain in response to specific stimuli or events, and they reflect the neural processing associated with those stimuli. Besides, $I/P$ represent inputs, $O/P$ for output, and $W$ represent Weights. Recently, [77] proposes a neural network model based on Bi-LSTM architecture for classifying hearing loss types using tonal audiometry data. The model achieves 99.33% classification accuracy on external datasets. The system can assist general practitioners in independently classifying audiometry results, reducing the burden on audiologists and improving diagnostic accuracy. The study, that may help assisting CI patients, aims to surpass the current SOTA accuracy rate of 95.5% achieved through DT.
III-F AE-based methods
@@
As a point of the case, the scientific paper [18] delves into the pivotal objective of enhancing speech perception for CI users in noisy conditions, recognizing the critical role of noise reduction (NR) in this pursuit. The proposed method, named DDAE-NR, has been proven effective in restoring clean speech. The study focuses on evaluating the DDAE-based NR using envelope-based vocoded speech, mimicking CI devices. The procedure of DDAE-based NR can be split into two main stages: training and testing. During the training phase, a collection of pairs of noisy and clean speech signals is prepared. These signals are initially transformed into the frequency domain using an fast Fourier transform (FFT). The logarithmic amplitudes of the noisy and clean speech spectra are then used as inputs and outputs, respectively, for the DDAE model.
Key findings underscore the superior intelligibility of DDAE-based NR in vocoded speech compared to SOTA conventional methods, indicating its potential implementation in CI speech processors. However, the study acknowledges the use of noise-vocoded speech simulation for evaluation and emphasizes the need for further validation with real CI recipients in clinical settings, addressing potential inconsistencies in the transition to actual CI devices.
@@
In addition, the study presented in [85] aims to enhance the accuracy and robustness of intra cochlear anatomy (ICA) segmentation, a vital component in preoperative decisions, insertion planning, and postoperative adjustments for CI procedures. The ICA includes structures such as scala tympani (ST), scala vestibul (SV), and the active region (AR). The researchers employed two segmentation methods, active shape model (ASM) and DL based on 3D U-Net AE, and combined them to achieve improved accuracy and robustness. A two-level training strategy involved pretraining on clinical CTs using ASM and fine-tuning on specimens’ CTs with ground truth. Results demonstrated that DL methods outperformed ASM in accuracy. While a trade-off between accuracy and robustness was observed, the combined DL and ASM approach showed improvements in both aspects. The study concludes that the proposed DL and ASM method effectively balances accuracy and robustness for ICA segmentation, highlighting the potential of DL-based methods, especially when integrated with ASM, to enhance CI procedures.
The proposed min-max similarity (MMS) methodology in [69] represents a groundbreaking approach to semi-supervised segmentation networks, particularly in the context of medical applications such as endoscopy surgical tool segmentation and CI surgery. MMS is introduced through dual-view training with contrastive learning, utilizing classifiers and projectors to create negative, positive, and negative pairs. The inclusion of pixel-wise contrastive loss ensures the consistency of unlabeled predictions. In the evaluation phase, MMS was tested on four public endoscopy surgical tool segmentation datasets and a manually annotated CI surgery dataset. The results demonstrate its superiority over SOTA semi-supervised and fully supervised segmentation algorithms, both quantitatively and qualitatively. Notably, MMS exhibited successful recognition of unknown surgical tools, providing reliable predictions, and achieved real-time video segmentation with an impressive inference speed of about 40 frames per second. This signifies the potential of MMS as a highly effective and efficient tool in medical image segmentation, showcasing its applicability in real-world surgical scenarios.
@@ III-G RL-based methods
IV Applications of DL-based medical CI
This section explores the application of deep learning in the field of cochlear implants, encompassing tasks such as speech denoising and enhancement, segmentation for precise identification and analysis of cochlear structures, thresholding, imaging, localization of CI, and more. Figure 12 provides a comprehensive overview of AI-based applications for CIs and theirs associated benefits. Furthermore, Table V summarizes various applications based on AI techniques, highlighting their performance, and pros and cons.
@@ IV Applications of DL-based medical CI
IV-A Speech denoising and enhancement
@@
TABLE V: Summary of the performance and limitations of specific DL applications dedicated to CIs. In cases where multiple tests are conducted, only the best performance is reported.
Ref. Year DLM Description BP Limitations PLA
[93] 2021 NC and DDAE_T The DL model comprises Siren noise at 6dB, a classifier, and the DDAE. The transfer learning (TL) is incorporated to help reduce the number of parameters in the model. PESQ= 3.037
@@
The performance of the DL model can be degraded when there is a mismatch between the training and testing data. No
[24] 2023 DeepACE Takes the raw audio signal as input and generates a denoised electrodogram, which represents the electrical stimulation patterns applied to the electrodes over time. STOI =0.807 Heavy data reliance impedes performance with insufficient or biased data, posing generalization challenges. Computational complexity, interpretability issues, and vulnerability are concerns. Yes [^17] https://github.com/APGDHZ/DeepACE2.0
[31] 2021 U-Net The proposed approach involves using deep virtual sensing to estimate microphone signals, improving speech quality in cocktail party scenarios for hearing aid and CI users. PESQ=2, STOI=0.53, SI-SDR=-28.91dB Challenges include long delay impracticality for hearing aids, binaural cues omission, insufficient network optimization, and exploring robustness in high-reverberation environments. No
[94] 2022 CNN A web-based research platform called Nautilus, which utilizes DL techniques for automated image processing in cochlear implantation-related studies. DCS= 80 ±$\pm 3$%, ASD= 0.19 ±$\pm 0.04$ mm Dependence on large amounts of annotated training data poses a challenge. Ensuring generalizability across various imaging conditions and populations is crucial. Yes181818nautilus_info@oticonmedical.com available upon reasonable request
[68] 2023 AE+Res-CA Is a vestibule segmentation network for CT images. It is based on the basic encoder-decoder framework and incorporates Res-CA Block, GCPFE and ACE-Loss Function DCS=94.77 ±$\pm 2.45$%, ASD=0.06 mm Limited training data, small object handling, generalization capability, and computational resource constraints are areas of concern. No
[95] 2019 cGAN Is a DL-based method for accurately localizing CI electrode contacts in CT images, facilitating customization of CI settings. success rate of localization =96.7 coarse resolution, metal artifacts, difficulty isolating contacts from cortical bones, site-dependent image quality, and variability among electrode array manufacturers. No
[34] 2022 DRL Is a novel method for accurately extracting ROIs in the inner ear from CT scans. The approach achieves high precision and demonstrates promising results for surgical planning in cochlear implantation. IoU= 0.84
@@
Small dataset, image quality variability, and computational requirements not specified. No
[86] 2020 U-Net Is a multi-scale DL framework for automatic segmentation and measurement of the human cochlea in clinical ultra-high-resolution CT images, with potential applications in personalized CI surgery. DSC=0.90 ±$\pm 0.03$
@@
The work include a small dataset, potential variability in image quality, lack of external validation, and limited assessment of clinical utility and computational requirements. No
[96] 2023 UNETR Explores the feasibility of using a DL method based on the UNETR model for automatic segmentation of the cochlea in temporal bone CT images. DSC=0.92 The small dataset used, the variability in image quality, and the absence of specifications regarding computational requirements. No
[97] 2019 3D U-NET Is a two-level training approach using a DL method to accurately segment the intra-cochlear anatomy in head CT scans. The method combines an active shape model-based method and a 3D U-Net model DSC=0.87 Limited dataset, variable image quality, lack of external validation, limited assessment of clinical utility, and no specifications on computational requirements. No
[36] 2023 AlexNet The study assesses repeatability, thresholds, and tonotopic patterns using a DL-based algorithm, providing insights into inner ear function and potential clinical applications. Acc=83.8% Potential dependence on the quality of the input data, limited generalizability to different patient populations or implant systems, and the need for further external validation and comparison with expert visual inspection. No
[…]
@@ IV-B Imaging
IV-C Segmentation
DL, ML, and AI have revolutionized CI segmentation, enabling precise identification and analysis of cochlear structures in various imaging modalities. Li et al. [96] applied a UNETR model to automatically segment cochlear structures in temporal bone CT images, enhancing surgical planning and cochlear implantation outcomes. Reda et al. [104] developed an automatic segmentation method for intra-cochlear anatomy in post-implantation CT scans, facilitating the customization of sound processing strategies for individual CI recipients. Moudgalya et al. [105] employed a modified V-Net CNN to segment cochlear compartments in $\mu$CT images, enabling precise quantification of local drug delivery for potential treatment of sensorineural hearing loss. Wang et al. [106] focused on metal artifact reduction and intra-cochlear anatomy segmentation in CT images using a multi-resolution multi-task deep network, benefiting CI recipients. Heutink et al. [86] developed a DL framework for the automatic segmentation and analysis of cochlear structures in ultra-high-resolution CT images, providing accurate measurements for personalized implant planning in cochlear surgery. Zhang et al. [97] utilized a 3D U-Net DL method to achieve accurate segmentation of intra-cochlear anatomy in head CT images, facilitating optimal programming of CIs and improving hearing outcomes. These studies highlight the significant impact of DL, ML, and AI in advancing CI segmentation, ultimately leading to improved patient care and treatment outcomes. Recently, Zhu et al. [107] proposes an uncertainty-aware dual-stream network, called UADSN, for facial nerve segmentation in CT scans for cochlear implantation surgery. UADSN combines 2D and 3D segmentation streams and uses consistency loss to improve accuracy in uncertain regions. The network achieves superior performance compared to other methods on a facial nerve dataset, with an emphasis on topology preservation.
IV-D Thresholding
@@
IV-E Localization of CI
DL methods have been instrumental in CI localization applications, providing accurate and automated solutions. Chi et al. [95] proposed a DL-based method for precise localization of electrode contacts in CT images. Their approach utilized cGANs to generate likelihood maps, which were then processed to estimate the exact location of each contact. Radutoiu et al. [34] focused on the automatic extraction of ROIs in full head CT scans of the inner ear. By leveraging AI, they achieved high precision in ROI localization, facilitating accurate surgical planning for insertion. Noble et al. [110] and Zhao et al. [111, 112], developed AI-based systems to automatically identify and position electrode arrays in CT images. These technologies enable large-scale analyses of the relationship between electrode placement and hearing outcomes, leading to potential advancements in implant design and surgical techniques. Heutink et al. [86] employed DL for the automatic segmentation and localization of the cochlea in ultra-high-resolution CT images. This approach allows for precise measurements that can be used in personalized planning, reducing the risk of intra-cochlear trauma and optimizing surgical outcomes. These studies showcase the significant contributions of DL and AI in localization applications, enabling accurate and efficient identification, positioning, and analysis of electrode arrays and facilitating improved surgical planning and outcomes. Burkart et al. [113] investigates the influence of sound source position and electrode placement on the stimulation patterns of CI under noise conditions. The study utilizes a measurement setup to simulate realistic listening scenarios. The results reveal that the effectiveness of CI noise reduction systems is influenced by these factors, and artificial intelligence fitting algorithms should be considered to optimize CI performance.
IV-F Other